OpenStudy (anonymous):

(1/32)^5x-1 =4^3x-1 solve for x

OpenStudy (anonymous):

I think you are missing some parentheses

OpenStudy (anonymous):

my teacher wrote it that way

OpenStudy (anonymous):

(1/32) to the 5x-1 = 4 to the 3x-1

OpenStudy (anonymous):

Take the log base (1/32) to get: \[\log_{\frac{1}{32}}(4^{3x-1}) = 5x-1 \] remembering that you can bring exponents out of logs you get: $$(3x-1)log_{\frac{1}{32}}(4) = 5x-1$$ now to get rid of the log base (1/32), you apply the rule $$log_n(m) = \frac{ln(m)}{ln(n)}$$ you get: $$\frac{ln(4)}{ln(1/32)} = \frac{5x-1}{3x-1}$$ and then doing some simple algebra and solving for x you get x = 7/31

OpenStudy (anonymous):

woohoo thats the answer i got!!